a. How many naturally occurring isotopes does magnesium have?

b. What is the weighted-average atomic mass of Magnesium in a sample that has the following isotopic abundance: 80%, 90%,99% (calculate based on the smallest to greatest values of the masses of the isotopes with the percent order, respectively)?
Answer #1

Please find attached the solution to the above question.
Answer #2

a. Naturally Occurring Isotopes of Magnesium

Magnesium has 3 naturally occurring isotopes:

  • ²⁴Mg (abundance ~79%),

  • ²⁵Mg (abundance ~10%),

  • ²⁶Mg (abundance ~11%).

Answer: 3



b. Weighted-Average Atomic Mass of Magnesium

Given:

  • Isotopic abundances: 80%90%99% (ordered from smallest to greatest isotope mass).

  • Assume isotope masses (in atomic mass units, u) follow natural order:

    • Lightest isotope (m1) = 24 u (²⁴Mg),

    • Middle isotope (m2) = 25 u (²⁵Mg),

    • Heaviest isotope (m3) = 26 u (²⁶Mg).


Calculation:

  1. Normalize abundances (must sum to 100%):
    Given percentages are 80%90%99% → Total = 80+90+99=269%.
    Normalized abundances:

    Abundance1=8026929.74%,Abundance2=9026933.46%,Abundance3=9926936.80%.

  2. Weighted-average atomic mass:

    Average mass=(0.2974×24)+(0.3346×25)+(0.3680×26).=7.1376+8.365+9.568=25.0706u.

Answer: 25.07u (rounded to 2 decimal places).


answered by: anonymous